OpenStudy (anonymous):

an estate was left to four children one got 1/5, second got 2/5, third got 1/10 how much did the fourth one get? i need help working it out.

OpenStudy (anonymous):

I know the lcd is what i have to find and would that be 10?

OpenStudy (anonymous):

what is the total estate? wen you say 1/5 .. you mean 1/5th of the total?

OpenStudy (anonymous):

3/10

OpenStudy (anonymous):

the rectangular prism has a length of 3.0cm a width of 2.2cm and a height of 7.5cm what is the surface area in square centimeters

OpenStudy (anonymous):

let x be the total estate: fourth share=x-x(1/5+2/5+1/10)=3/10x

OpenStudy (anonymous):

the rectangular prism has a length of 3.0cm a width of 2.2cm and a height of 7.5cm what is the surface area in square centimeters